OpenStudy (anonymous):

help me solve this ? 5x + 17 ≥ 7

OpenStudy (anonymous):

5x - 17 = -x + 7 Add both sides by x: 5x + x - 17 = -x + x + 7 6x - 17 = 7 Add both sides by 17: 6x - 17 + 17 = 7 + 17 6x = 24 Finally, divide both sides by 6: 6x/6 = 24/6 x = 24/6 x = 4 I hope this helps!

OpenStudy (haseeb96):

\[5xge7-17\] \[5x \ge-10\]\[x \ge-2\] this is the solution of it and answer is also correct